did i solve this correctly? 5x+2y=7x-14, step one 5x+2y-5x = 7x-14-5x, step two 2y = 2x-14, step three 2y/2 = 2x-14/2, step four y=x-7 as answer?

Answer by Alan3354(69443)

5x+2y=7x-14
-2x + 2y = -14
2y = 2x - 14
y = x-7
Looks good to me.